Splicing has completed and most services are back online now. Please power cycle your modem if service is not restored and contact support if issues persist.

Engineers have advised due to unforeseen conditions the splicing has been delayed. The revised ETA is for 22:15 AEST. The next update will be shared by 23:00 AEST.

Field engineers have confirmed that cable has arrived and hauling has been completed. Fibre technicians are in a process of preparing for splicing with an ETA of 20:45 AEST. The next update will be shared by 21:45 AEST.

Engineers have found the exact location of the break under the driveway at a shopping centre. The fiber technician has successfully found a solution to run the cable under the driveway. The replacement cable is on route, The expected ETA to arrive at the site is 120mins, due to peak hour traffic. Preparation work to haul the cable will be looked at whilst waiting for the cable to arrive. Once the cable has arrived Hauling process will begin, cable length to hauled is approximately 125m. ETA for splicing will be informed as soon hauling starts.

The next update will be shared by 19:30 AEST


Engineers have identified a fiber break in Sherif Road, Lonsdale.

Field operations and fibre team have been engaged with an ETA of 30mins. The next update will be shared by 17:30 AEST.


We have detected an issue affecting some NBN customers in South Australia. Affected areas include MORPHETT VALE, HALLETT COVE, WOODCROFT, LONSDALE and surrounding suburbs . Affected customers will be offline.

Began at:

Affected components
  • NBN
    • NBN SA